OpenStudy (anonymous):

What is the x-intercept of the line with the equation ? 8x-1 over 3y=16 ( , 0)

OpenStudy (anonymous):

@otakukid99

OpenStudy (anonymous):

@Ontario_C

OpenStudy (anonymous):

@otakukid99

OpenStudy (anonymous):

...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can u help

OpenStudy (anx):

x intercepts are found by setting y equal to zero set up as 8x = 16 +1/3y set y equal to zero 8x = 16 + (1/3)(0) simplify 8 x = 16 divide both sides by 8 x = 2

OpenStudy (anx):

Took me a minute to figure out how the equation was set up, lol

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

@anx can u help wth a few more

OpenStudy (anx):

Sure, watcha got